OpenStudy (anonymous):

HELPPP!!! Find the sum of the even numbers from 1 to 200 inclusive.

OpenStudy (anonymous):

does anyone know??

OpenStudy (anonymous):

=2 + 4 + ... + 200 =2(1+2+3+...+100) =2*((101*100)/2) = 101*100 =10100

OpenStudy (anonymous):

im not doubting u but are u positive on this answer @jeanlouie

OpenStudy (helder_edwin):

don't doubt him/her. 10100 is correct.

OpenStudy (anonymous):

ok thank you @helder_edwin

OpenStudy (anonymous):

thank you also @jeanlouie

OpenStudy (anonymous):

hope that help :)

OpenStudy (anonymous):

can either of you answer this one?>>>sin θ = -3/5, find sin 2θ given that θ is in the 3rd quadrant.

OpenStudy (anonymous):

θ is in the 3rd quadrant then cosθ < 0. (sinθ)^2 + (cosθ)^2 = 1 => (cosθ)^2= 1-(-3/5)^2 = 16/25 => cosθ = -4/5 So, sin2θ = 2sinθcosθ = 2 * (-3/5) * (-4/5) = 24/25

OpenStudy (anonymous):

do u think u could help me with some more but um going to close this question and open a new one and thank u